Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $160,000 per year. Its operating results for last year were as follows:

Sales $ 2,080,000
Variable expenses 1,040,000
Contribution margin 1,040,000
Fixed expenses 160,000
Net operating income $ 880,000

5. The sales manager is convinced that a 14% reduction in the selling price, combined with a $64,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. Do you recommend implementing the sales manager's suggestions?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.40 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $880,000 net operating income as last year?
Answer #1

Given data:

Total sales value = $ 2,080,000

Selling price per unit = $ 80

No of units sales = 2,080,000/80 =26,000 units

Variable expenses per unit = $ 40

Total variable exp = 26,000*40 = $ 1,040,000

Total fixed cost = $ 160,000

Net operating income = $ 880,000

5. If Selling price reduced by 25% per unit, combined with $ 64,000 increase in advertising exp, would increase unit sales by 25%:

a) Calculating net operating income:

New selling price per unit = ($80- 80*25/100) =$ 60

New total selling unit = (26,000+26,000*25/100) =32,500 units

Total sales =32,500*60 = $ 1,950,000

Variable cost per unit = $ 40

Total variable cost =32,500*40 = $ 1,300,000

New fixed cost = $ 160,000+ 64,000 = $234,000

Contribution = sales – variable cost =1,950,000-1,300,000 =$ 650,000

Contribution ratio = contribution/ sales*100

So, contribution ratio = 650,000/1,950,000*100 =33.33%

Sales= fixed cost+net operating income /contribution ratio

1,950,000 = 234,000+ net operating income /33.33%

1,950,000*33.33/100 =234,000+net operating income

650,000-234,000 = net operating income

Net operating income = $ 416,000

Proof:

Sales =                         $ 1,950,000

Variable exp =              $ 1,300,000

Contribution margin = $    650,000

Fixed exp = $     234,000

Net operating income= $   416,000

b) sales manager suggestion is not recommendable to implement, because the new total net operating income is lesser than old. I,e., new operating income = $ 416,000, but old net operating income $880,000.

6) Increase sales commission per unit $ 2.40, combined with some increase in advertising exp, would increase sales by 25%, how much increase the advertising exp to earn a $ 880,000 as net operating income:

Selling price per unit = $ 80

New total selling unit =26,000+26,000*25/100 = 32,500 units

Total sales = 32,500*80 = $ 2,600,000

Variable exp = $ 40+ 2.40 = $ 42.40

Total variable exp = 32,500*42.40 = $ 1,378,000

Contribution = Sales-variable exp =2,600,000-1,378,000 =$ 1,222,000

Contribution ratio = contribution / sales*100 = 1,222,000/2,600,000*100 = 47%

Sales =total new fixed cost+ net operating income/contribution ratio

Total new fixed cost+net operating income =sales*contribution ratio

Total new fixed cost + 880,000 = 2,600,000*47/100

Total new fixed cost + 880,000 = 1,222,000

Total new fixed cost = 1,222,000-,880,000 =342,000

Therefore advertising exp =total new fixed cost – old fixed cost

342,000- 160,000 = $ 182,000
Answer #2
Req 5-aContribution Income   statement


 Last Year Proposed


                 26,000 Units               32,500 Units 


Total Per Unit Total Per Unit 

Sales           2,080,000                         80.00        2,236,000                 68.80

Less: Variable expense           1,040,000                         40.00        1,300,000                 40.00

Contribution margin           1,040,000                         40.00            936,000                 28.80

Fixed expenses              160,000
            224,000

Net Income              880,000
            712,000






Req 5-bNet Operating Income Decrease by            (168,000)








Req 6The amount by which advertising can be increased              182,000









Working Note :




Contribution margin per unit (Last year)                   40.00



Less: increase in sales comm                     (2.40)



Prposed contribution margin                   37.60



Multiplied by proposed sales after increase                 32,500



Proposed total contribution margin           1,222,000



Less: Fixed expenses            (160,000)



Less: Target Income            (880,000)



The amount by which advertising can be increased              182,000



source: managerial accounting
answered by: anonymous